A triangle has a perimeter of 18 inches. If one side has length 8 inches, find the other two sides if their lengths are in the ratio of 2units : 3units.
Answer by cleomenius(959)

18 = 8 + 2x + 3x.
10 = 5x
x = 2
18 = 8 + 2(2) + 3(2)
18 = 8 + 4 + 6
Cleomenius.